sam (.sam.):

In an AP, d=1.6, a=6 \[S_{n}=pn^2+qn \] Find p and q

sam (.sam.):

@experimentX @lgbasallote @Callisto Try this? :)

OpenStudy (callisto):

I'm going to give wrong answers :( T(n) = 6+1.6(n-1) S(n) = [6+ ( 6+1.6(n-1)) ] n/2 = (12+1.6n -1.6) n/2 = (1.6n^2 + 10.4n)/2 = 0.8n^2 + 5.2n p= 0.8, q=5.2

OpenStudy (lgbasallote):

a is first term right? A_n = 6 +(n-1)1.6 A_n = 6 + 1.6n - 1.6 A_n = 4.4 + 1.6n S_n = n(A_1 + A_n)/2 pn^2 + qn = n(6 + 4.4 + 1.6n)/2 n(pn + q) = n(10.4 + 1.6n)/2 pn + q = 5.2 + 0.8n

OpenStudy (lgbasallote):

lol i didnt know what to do with that...genius caLLISTO!

sam (.sam.):

@Callisto is correct :D

OpenStudy (callisto):

Oh... I must be dreaming :(....